OpenStudy (anonymous):

What is the exact area of a circular courtyard with a radius of 22 ft? A. 1936π ft² B. 484π ft² C. 44π ft² D. 22π ft²

jhonyy9 (jhonyy9):

do you know formula for area of a circle ?

OpenStudy (asevilla5):

a=π•r^2 this is the formula

jhonyy9 (jhonyy9):

yes so just substitute the added dates for r and pi and will get the right answer sure easy

OpenStudy (asevilla5):

first plug in the in the 22 for r and you should get the answer @CUTIEGIRL01 very easy

OpenStudy (anonymous):

is the answer is b 484

OpenStudy (anonymous):

@Asevilla5

OpenStudy (asevilla5):

correct @CUTIEGIRL01

OpenStudy (anonymous):

Which expression can be used to find the area of a circle with a diameter of 40 cm? A. π · 20 B. π · 40 C. π · 202 D. π · 402

jhonyy9 (jhonyy9):

so than the diameter is 40 how many will be the radius ?

OpenStudy (asevilla5):

area of a circle is π * r^2 The radius is half the diameter so the radius in this case is 20 cm. So r = 20 π * 20^2 = π · 402

OpenStudy (anonymous):

the answer is 402

OpenStudy (asevilla5):

good job

OpenStudy (anonymous):

40 square

OpenStudy (anonymous):

the 2 is the squared sign on c and d so it is either a 20 b 40 c 20 squared or c 40 squared

OpenStudy (anonymous):

@Asevilla5

jhonyy9 (jhonyy9):

how you have got 402 above like result for area ?

jhonyy9 (jhonyy9):

than the radius is 20 what is 40/2

OpenStudy (anonymous):

is the answer 20

jhonyy9 (jhonyy9):

20 is just the lenght of radius

jhonyy9 (jhonyy9):

so this mean pir^2 =pi*20^2

jhonyy9 (jhonyy9):

or not ?

OpenStudy (anonymous):

so the answer is 40

jhonyy9 (jhonyy9):

why ? 20^2 = ?

OpenStudy (anonymous):

what does the ^ sign mean

jhonyy9 (jhonyy9):

exponent 20 on power 2

OpenStudy (anonymous):

so how do i find the answer

jhonyy9 (jhonyy9):

20^2 =400

OpenStudy (anonymous):

but that is not a choice

OpenStudy (triciaal):

A = pi * r ^2 r = diameter/2 = 40/2 = 20 leave pi as pi A = pi * 20 ^2 option c

OpenStudy (anonymous):

thank you all for your help